# Approximate predictive control barrier function for discrete-time systems

### v1

@@ 4.2 Inherent robustness
The system $x(k+1)\in F(x(k),w(k))$ is input-to-state stable in $\mathcal{D}$ with respect to the compact, non-empty and 00-input invariant set [^2] A set $\mathcal{S}\subset\mathcal{D}$ is denoted 00-input invariant for the dynamics $x(k+1)\in F(x(k),w(k))$ if it is invariant for the dynamics $x(k+1)\in F(x(k),0)$. $\mathcal{S}$ if there exists a $\mathcal{KL}$ function [^3] According to [8, Definition B.3], a function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{N}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ belongs to class $\mathcal{KL}$ if it is continuous and if, for each $t\geq 0$, $\beta(\cdot,t)$ is a class $\mathcal{K}$ function and for each $s\geq 0$, $\beta(s,\cdot)$ is non-increasing and satisfies $\lim_{t\rightarrow\infty}\beta(s,t)=0$. $\beta$ and a $\mathcal{K}$ function $\sigma$ such that, for each $x\in\mathcal{D}$, with $\mathcal{D}$ compact and robustly positively invariant, and each disturbance sequence $\mathbf{w}=(w(0),w(1),\dots)$ in $\ell_{\infty}$
@@
for all $i\in\mathbb{N}_{\geq 0}$, where $\phi(i;x(0),\mathbf{w}_{i})$ is any solution of (16) at time $i$, if the initial state is $x(0)$ and the input sequence is $\mathbf{w}_{i}:=(w(0),w(1),\dots,w(i-1))$.
We note that with regards to input-to-state stability, the disturbances $w(k)$ are considered as the input of the system. Following the definition of ISS, a corresponding ISS Lyapunov function can be defined, which we denote as ISS control barrier function in this paper to distinguish stability of singletons from stability of sets.
@@
Let $\mathcal{S}$ and $\mathcal{D}$ be compact and non-empty, with $\mathcal{S}$ 00-input invariant, $\mathcal{D}$ robustly positively invariant and $\mathcal{S}\subset\mathcal{D}$. A function $h:\mathcal{D}\rightarrow\mathbb{R}_{\geq 0}$ is an ISS-CBF for system $x(k+1)\in F(x(k),w(k))$ if there exist class $\mathcal{K}$ functions $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ and $\sigma$ such that for all $x\in\mathcal{D},w\in\mathbb{R}^{p}$,
@@
By leveraging the fact that the PCBF defined in (10) is a CBF according to Definition 4 according to Corollary 4.8, we are able to relate it to the notion of ISS-CBF in Definition 4.11.
@@
Let Assumptions 1 and 2 hold. The closed-loop system (16) is ISS according to Definition 4.10.
Proof 4.15.
We show ISS by proving that $h_{\mathrm{PB}}$, as defined in (10), is an ISS-CBF according to Definition 4.11. From Corollary 4.8, it follows that $h_{\mathrm{PB}}$ is a CBF according to Definition 4. It therefore holds that $\mathcal{S}_{\mathrm{PB}}$ is 00-input invariant for the difference inclusion $x(k+1)\in F(x(k),0)$ as it holds that $u\in\mathcal{U}_{h_{\mathrm{PB}}}$ according to (16). Furthermore, from Definition 4, it follows that there exist $\alpha_{1}$ and $\alpha_{2}$ such that $\alpha_{1}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})\leq h_{\mathrm{PB}}(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})$. It therefore remains to be shown that $\sup_{x^{+}\in F(x,w)}h_{\mathrm{PB}}(x^{+})-h_{\mathrm{PB}}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+\sigma(\|w\|)$. Given the definition of $\mathcal{U}_{h_{\mathrm{PB}}}$, it holds that $\forall x\in\mathcal{D}_{\mathrm{PB}}$, we have
$\sup_{x^{+}\in F(x,w)}h_{\mathrm{PB}}(x^{+})=\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,w))$
From the Heine-Cantor theorem in [39, Theorem 4.19], it follows that the continuous functions $h_{\mathrm{PB}}$ and $f$ are uniformly continuous on the compact set $\mathcal{D}_{\mathrm{PB}}$. It therefore holds that there exist class $\mathcal{K}$ functions $\sigma_{h_{\mathrm{PB}}}$ and $\sigma_{f}$ according to [40, Lemma 1] such that
$\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,w))$
$\leq$ $\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma_{h_{\mathrm{PB}}}\circ\sigma_{f}(\|w\|)$
$=$ $\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma(\|w\|),$
where $\sigma=\sigma_{h_{\mathrm{PB}}}\circ\sigma_{f}$ is also a class $\mathcal{K}$ function. By using $\mathcal{U}_{h_{\mathrm{PB}}}=\{u\in\mathcal{U}\mid h_{\mathrm{PB}}(\tilde{f}(x,u,0))-h_{\mathrm{PB}}(x)\leq-\Delta h_{\mathrm{PB}}(x)\}$ for $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, we have
$\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma(\|w\|)$
$\leq$ $h_{\mathrm{PB}}(x)-\Delta h_{\mathrm{PB}}(x)+\sigma(\|w\|)$
$\leq$ $h_{\mathrm{PB}}(x)-\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+\sigma(\|w\|),$
where $\alpha_{3}$ exists according to Lemma 1. For any $x\in\mathcal{S}_{\mathrm{PB}}$, we have $\mathcal{U}_{h_{\mathrm{PB}}}=\{u\in\mathcal{U}\mid h_{\mathrm{PB}}(\tilde{f}(x,u,0))\leq 0\}$, resulting in
$\sup_{u\in\mathcal{U}_{h_{\mathrm{PB}}}}h_{\mathrm{PB}}(\tilde{f}(x,u,0))+\sigma(\|w\|)$
$\leq\sigma(\|w\|)$
$=h_{\mathrm{PB}}(x)-\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+\sigma(\|w\|),$
as $h_{\mathrm{PB}}(x)=\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})=0$ for $x\in\mathcal{S}_{\mathrm{PB}}$. Therefore, for any $x\in\mathcal{D}_{\mathrm{PB}}$ the decrease condition (17b) holds and $h_{\mathrm{PB}}$ is an ISS-CBF according to Definition 4.11 and the closed-loop system (16) is ISS, which follows from Lemma 4.12.
Given that the PCBF defines an ISS-CBF for system (16), it holds that the nominal predictive control problem (10) admits inherent robustness properties. In other words, it holds that for any state $x(0)\in\mathcal{D}_{\mathrm{PB}}$, a neighbourhood of $\mathcal{S}_{\mathrm{PB}}$ is stable for the system (16).
5 EXPLICIT APPROXIMATION OF THE PCBF
In this section, we present the main proposed approximation-based algorithm. Given the theoretical foundation established in Section 4, it holds that a closed-loop system (16) is inherently robust with respect to exogeneous disturbances and in this section, we show how possible approximation errors can be treated in a similar manner. Furthermore, we propose a continuous extension for the PCBF, which is negative whenever $h_{\mathrm{PB}}$ is 00, such that safety filter interventions due to approximation errors within $\mathcal{S}_{\mathrm{PB}}$ are reduced.
5.1 Approximate PCBF
The optimal value function in (10) is a CBF according to Definition 4 and encodes the information of the entire prediction horizon of the predictive optimisation problem with respect to potential constraint violations. Therefore, a time horizon independent optimization problem that could be solved instead of (10) is given by
$\min_{u\in\mathcal{U}}h_{\mathrm{PB}}(f(x(k),u))$ (18)
Clearly, such a formulation benefits from the same guarantees as solving (10). Given a state $x\in\mathcal{S}_{\mathrm{PB}}$, any input $u\in\arg\min h_{\mathrm{PB}}(f(x(k),u))$ ensures that at the next time step, the system remains in the safe set $\mathcal{S}_{\mathrm{PB}}$, rendering it invariant. Similarly, for states in $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, by minimising the PCBF value, the decrease $\Delta h_{\mathrm{PB}}$ in (3a) is achieved and stability of $\mathcal{S}_{\mathrm{PB}}$ is ensured.
However, an explicit formulation of $h_{\mathrm{PB}}$ is generally not available. We therefore propose to approximate the function $h_{\mathrm{PB}}$ with a continuous regressor $\hat{h}_{\mathrm{PB}}$. Due to the continuity property of the PCBF in (10) and the fact that it can be sampled offline, the function is suitable for approximation, such that a computationally efficient algorithm can be leveraged. As the approximation $\hat{h}_{\mathrm{PB}}$ will be embedded in an optimisation problem, we consider neural networks in this paper due to the parametric nature of the regressor, allowing for a formulation independent of the number of training samples, compared to, e.g., kernel-based methods. Note that while in principle a safety filter policy $\pi(x,u_{\mathrm{p}})$ could be learned, this policy is generally not only non-continuous and multi-valued, but also has a higher input dimensionality for the data collection and training as $\pi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$. As an illustrative example, consider a car approaching an obstacle. If the performance input is given to steer straight towards the obstacle, an optimal solution for the safety filter problem is given by either steering fully to the left or to the right.
In order to train the regressor for the PCBF approximation, a dataset of samples has to be collected offline in the form of $\mathbb{D}=\{(x_{i},h_{\mathrm{PB}}(x_{i}))\}_{i=0}^{n_{d}}$. This data set can be sampled from the entire region of attraction $\mathcal{D}_{\mathrm{PB}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)\leq\gamma_{f}\alpha_{f}\}$, where $\gamma_{f}$ is the sublevel set of the domain of the terminal CBF $\mathcal{D}_{f}$ and $\alpha_{f}$ is the terminal penalty in (10a), or from a subset deemed relevant to the specific application. In the Appendix, we propose a geometric sampling method based on [41], which allows efficiently sampling the PCBF by performing a pseudo random walk over a preselected domain, using previous solutions as warmstarts to the next computed optimisation problem.
Given a trained explicit approximation $\hat{h}_{\mathrm{PB}}$ of $h_{\mathrm{PB}}$, the optimisation problem (18) could directly be modified to the optimisation problem $\min_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(f(x(k),u))$. Such an approach can be useful as an efficient control mechanism for convergence of states $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$ to $\mathcal{S}_{\mathrm{PB}}$ with no approximation errors, however, it does not easily allow for considering the primary objective considered in this paper, i.e. filtering inputs with respect to safety regards. Solving a subsequent safety filter problem (9) would still require the full information of the optimal slack variables across the prediction horizon. Instead of minimising the approximate PCBF value, the corresponding CBF decrease condition (3a) can directly be used as a constraint as is standard in CBF-based safety filters, see, e.g. [5] for continuous-time CBF applications. The optimisation problem which is solved at every time step is then given by
$\begin{split}\min_{u}&\|u-u_{\mathrm{p}}(k)\|\\ \textup{s.t. }&\hat{h}_{\mathrm{PB}}(f(x(k),u))\!-\!\hat{h}_{\mathrm{PB}}(x(k))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x(k)),\\ &u\in\mathcal{U},\end{split}$ (19)
where $c_{\Delta}\in(0,1]$ is a user defined hyperparameter and the function $\Delta\hat{h}_{\mathrm{PB}}(x)$ is a decrease function of the approximate PCBF $\hat{h}_{\mathrm{PB}}(x)$. The main advantage of the proposed optimisation problem compared to the original PCBF mechanism, which solves (10) and (9) consecutively consists of the fact that the formulation in (19) is independent of the prediction horizon of the predictive control based optimisation problems. Instead, only a single decrease constraint needs to be fulfilled along with the input constraints and the only optimisation variable is $u\in\mathbb{R}^{m}$.
Given that the decrease function of $h_{\mathrm{PB}}(x)$, i.e. $\Delta h_{\mathrm{PB}}(x)$, is not explicitly available in a similar fashion to $h_{\mathrm{PB}}(x)$ itself, we propose an optimisation-based formulation for $\Delta\hat{h}_{\mathrm{PB}}$. The approximate decrease which is used in (19) is defined as
$\Delta\hat{h}_{\mathrm{PB}}(x):=\max_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u))$ (20)
From this formulation of the approximate decrease function, it follows that the hyperparameter $c_{\Delta}$ allows tuning the desired decrease rate of the closed-loop system for any $x\in\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$. By selecting $c_{\Delta}=1$, at every time step, the approximate PCBF $\hat{h}_{\mathrm{PB}}$ undergoes the maximal possible decrease, while choosing $c_{\Delta}\in(0,1)$ reduces the safety interventions of the proposed safety filter by reducing the amount that the approximate PCBF is required to decrease. This provides an additional degree of freedom compared to the PCBF mechanism, where at every time steps the minimal slacks in terms of the objective function (10a) are required to be used in the safety filter and relaxing the corresponding decrease is a theoretically challenging problem.
@@ 5.1 Approximate PCBF
An additional advantage of the explicit approximation of the PCBF is given by the fact that proposed inputs $u_{\mathrm{p}}(k)$ can be very efficiently certified in terms of satisfying optimality with respect to (19) if $u_{\mathrm{p}}(k)\in\mathcal{U}$. It suffices to perform a simple evaluation, whether $\hat{h}_{\mathrm{PB}}(f(x(k),u_{\mathrm{p}}(k)))\leq 0$, as it then holds that $u_{\mathrm{p}}(k)$ is an optimal solution to (19). This function evaluation can thereby reduce the need for solving two optimisation problems online at every time step. The main proposed algorithm is then given in Algorithm 1, where first, the proposed input is certified and if the input is not deemed to be safe, the maximal decrease in (20) is computed and then used in the safety filter (19).
Algorithm 1 Approximate PCBF safety filter
1:  for $k\geq 0$ do
2:     Measure state $x(k)$
3:     Obtain the proposed control input $u_{\mathrm{p}}(k)$
4:     if $u_{\mathrm{p}}(k)\in\mathcal{U}\wedge\hat{h}_{\mathrm{PB}}(f(x(k),u_{\mathrm{p}}(k)))\leq 0$ then
5:        Apply $u(k)\leftarrow u_{\mathrm{p}}(k)$
6:     else
7:        Compute $\Delta\hat{h}_{\mathrm{PB}}(x(k))$ using (20)
8:        Apply $u(k)\leftarrow u^{*}$, solution of (19)
9:     end if
10:  end for
Given the proposed algorithm, we can now provide theoretical guarantees of the closed-loop behaviour of the system despite possible approximation errors. First, we show that the optimisation problem (19) is feasible using the maximum decrease (20) for any $x\in\mathcal{D}_{\mathrm{PB}}$.
Proposition 5.17.
The optimisation problem (19), where $\Delta\hat{h}_{\mathrm{PB}}(x)$ is defined in (20) is feasible for any $x(k)\in\mathbb{R}^{n}$ and any $c_{\Delta}\in(0,1]$.
Proof 5.18.
The optimisation problem (20) is feasible for any $x\in\mathbb{R}^{n}$ by choosing any $u\in\mathcal{U}$. Given an optimal solution $u^{*}$ to (20), it holds that $u^{*}$ is feasible in (19) as $u^{*}\in\mathcal{U}$ and for any $x\in\mathbb{R}^{n}$, $\hat{h}_{\mathrm{PB}}(f(x,u^{*}))-\hat{h}_{\mathrm{PB}}(x)=-\max_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x)$ for all $c_{\Delta}\in(0,1]$.
In a second step, we can show that the closed-loop system is ISS with respect to the approximation errors of $\hat{h}_{\mathrm{PB}}$ if the approximation error is bounded according to the following assumption.
Assumption 3
Such a uniform bound for the approximation error always holds if the chosen regressor for $\hat{h}_{\mathrm{PB}}$ is continuous. Due to continuity of $\hat{h}_{\mathrm{PB}}$ and $h_{\mathrm{PB}}$ as well as compactness of $\mathcal{D}_{\mathrm{PB}}$, a maximum $\epsilon_{h}$ exists for $\max_{x\in\mathcal{D}_{\mathrm{PB}}}|h_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(x)|$ according to the extreme value theorem [8, Proposition A.7].
Theorem 5.19.
Let Assumptions 12 and 3 hold. Then, for small enough approximation error $\epsilon_{h}$, the closed-loop system (1) under application of Algorithm 1 is ISS according to Definition 4.10.
Proof 5.20.
We show that $h_{\mathrm{PB}}(x)$ is an ISS Lyapunov function according to Definition 4.11 for the closed-loop system $x(k+1)\in\hat{F}(x(k))=\{x^{+}\in\mathbb{R}^{n}\mid x^{+}=f(x(k),u),u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}\}$, where $\mathcal{U}_{\hat{h}_{\mathrm{PB}}}\coloneqq\{u\in\mathcal{U}\mid\hat{h}_{\mathrm{PB}}(f(x(k),u))\!-\!\hat{h}_{\mathrm{PB}}(x(k))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x(k))\}$. Given that inputs are selected as an optimal solution to (19), it holds that $u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}$ under application of Algorithm 1. As $h_{\mathrm{PB}}$ is a CBF according to Definition 4, the condition (17a) is satisfied. Furthermore, we can show that for any $u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}$, it holds that
$h_{\mathrm{PB}}(f(x,u))-h_{\mathrm{PB}}(x)$
$\leq$ $\hat{h}_{\mathrm{PB}}(f(x,u))-e_{h}(f(x,u))-\hat{h}_{\mathrm{PB}}(x)+e_{h}(x)$
$\leq$ $-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x)+2\epsilon_{h}$
$\leq$ $-c_{\Delta}\max_{u\in\mathcal{U}}\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u))+2\epsilon_{h}$
$\leq$ $-c_{\Delta}\Delta h(x)+4\epsilon_{h},$
where we used $c_{\Delta}\in(0,1]$ in the last inequality, such that (17b) is satisfied. The set $\mathcal{S}_{\mathrm{PB}}$ is 00-invariant with $\epsilon_{h}=0$ and the set $\mathcal{D}_{\mathrm{PB}}$ is robust invariant, if $\epsilon_{h}$ is sufficiently small, i.e., for all $x\in\mathcal{D}_{\mathrm{PB}}=\{x\in\mathbb{R}^{n}\mid h_{\mathrm{PB}}(x)\leq\alpha_{f}\gamma_{f}\}$, it holds that $h_{\mathrm{PB}}(f(x,u))\leq\alpha_{f}\gamma_{f}$. This holds if $\epsilon_{h}\leq\frac{1}{4}(\alpha_{f}\gamma_{f}-\bar{c})$ with $\bar{c}\coloneqq\max_{x\in\mathcal{D}_{\mathrm{PB}}}h_{\mathrm{PB}}(x)-c_{\Delta}\alpha_{3}\circ\alpha_{1}^{-1}(h_{\mathrm{PB}}(x))\leq\alpha_{f}\gamma_{f}$ as we have $h_{\mathrm{PB}}(x)\leq\alpha_{f}\gamma_{f}$ and
$h_{\mathrm{PB}}(x)-c_{\Delta}\alpha_{3}(\|x\|_{\mathcal{S}_{\mathrm{PB}}})+4\epsilon_{h}$
$\leq$ $h_{\mathrm{PB}}(x)-c_{\Delta}\alpha_{3}\circ\alpha_{1}^{-1}(h_{\mathrm{PB}}(x))+4\epsilon_{h}\stackrel{{\scriptstyle!}}{{\leq}}\alpha_{f}\gamma_{f}$
It therefore holds that using the approximation-based safety filter (19), the system converges to and remains in a neighbourhood of the safe set $\mathcal{S}_{\mathrm{PB}}$.
Remark 5.21.

### v2

@@ 5.2 ISS of the Approximate PCBF Algorithm
The system $x(k+1)\in F(x(k),w(k))$ is input-to-state stable in $\mathcal{D}$ with respect to the compact, non-empty and 0-input invariant set [^2] A set $\mathcal{S}\subset\mathcal{D}$ is denoted 0-input invariant for the dynamics $x(k+1)\in F(x(k),w(k))$ if it is invariant for the dynamics $x(k+1)\in F(x(k),0)$. $\mathcal{S}$ if there exists a $\mathcal{KL}$ function [^3] According to [8, Definition B.3], a function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{N}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ belongs to class $\mathcal{KL}$ if it is continuous and if, for each $t\geq 0$, $\beta(\cdot,t)$ is a class $\mathcal{K}$ function and for each $s\geq 0$, $\beta(s,\cdot)$ is non-increasing and satisfies $\lim_{t\rightarrow\infty}\beta(s,t)=0$. $\beta$ and a $\mathcal{K}$ function $\sigma$ such that, for each $x\in\mathcal{D}$, with $\mathcal{D}$ compact and robustly positively invariant, and each disturbance sequence $\mathbf{w}=(w(0),w(1),\dots)$ in $\ell_{\infty}$
@@
for all $i\in\mathbb{N}_{\geq 0}$, all solutions of the system $\phi(i;x(0),\mathbf{w}_{i})$ at time $i$ if the initial state is $x(0)$, and the input sequence is $\mathbf{w}_{i}:=(w(0),w(1),\dots,w(i-1))$.
We note that with regards to input-to-state stability, the disturbances $w(k)$ are considered as the input of the system. Following the definition of ISS, a corresponding ISS Lyapunov function can be defined, which we denote as ISS control barrier function in this paper to distinguish stability of singletons from stability of sets. We use the following definition to characterize an ISS-CBF, implying ISS of the closed-loop as shown in [38].
@@
Let $\mathcal{S}$ and $\mathcal{D}$ be compact and non-empty, with $\mathcal{S}$ 0-input invariant, $\mathcal{D}$ robustly positively invariant and $\mathcal{S}\subset\mathcal{D}$. A function $h:\mathcal{D}\rightarrow\mathbb{R}$ is an ISS-CBF for system $x(k+1)\in F(x(k),w(k))$ if there exist class $\mathcal{K}$ functions $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ and $\sigma$ such that for all $x\in\mathcal{D},w\in\mathbb{R}^{p}$,
@@
By showing the existence of an ISS-CBF implies that the closed-loop system under the proposed approximate PCBF algorithm is ISS according to [8, Definition B.42].
@@
Let Assumptions 12 and 3 hold. If $\mathcal{D}_{\mathrm{PB}}$ is robustly positively invariant for the approximation $\hat{h}_{\mathrm{PB}}$ and disturbances $w(k)$, system (1) under application of Algorithm 1 is ISS.
Proof 5.22.
We show that $h_{\mathrm{PB}}(x)$ is an ISS-CBF according to Definition 5.18 for the closed-loop system $x(k+1)\in\hat{F}(x(k),w(k))=\{x^{+}\in\mathbb{R}^{n}\mid x^{+}=f(x(k),u,w(k)),u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}(x(k))\}$, where we define $\mathcal{U}_{\hat{h}_{\mathrm{PB}}}(x(k))\coloneqq\{u\in\mathcal{U}\mid\hat{h}_{\mathrm{PB}}(f(x(k),u,0))\!-\!\hat{h}_{\mathrm{PB}}(x(k))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x(k))\}$. Given that inputs are selected as an optimal solution to (11), it holds that under application of Algorithm 1, $u(k)\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}(x(k))$. As $h_{\mathrm{PB}}$ is a CBF according to Definition 4, the condition (20a) is satisfied. Furthermore, we can show that for any $u\in\mathcal{U}_{\hat{h}_{\mathrm{PB}}}(x)$, it holds $\forall w\in\mathbb{R}^{p}$ that
$\sup_{x^{+}\in\hat{F}(x,w)}h_{\mathrm{PB}}(x^{+})-h_{\mathrm{PB}}(x)$
$\leq$ $\hat{h}_{\mathrm{PB}}(f(x,u,w))+|e_{h}(f(x,u,w))|-\hat{h}_{\mathrm{PB}}(x)+|e_{h}(x)|$
$\leq$ $\hat{h}_{\mathrm{PB}}(f(x,u,0))-\hat{h}_{\mathrm{PB}}(x)+\sigma_{h}(\|w\|)+|e_{h}(f(x,u,0))|$
$+\sigma_{e}(\|w\|)+|e_{h}(x)|$
$\leq$ $-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x)+2\epsilon_{h}+\sigma(\|w\|)$
$\leq$ $-c_{\Delta}\max_{u\in\mathcal{U}}\left(\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u,0))\right)+2\epsilon_{h}+\sigma(\|w\|)$
$\leq$ $-c_{\Delta}\Delta h(x)+4\epsilon_{h}+\sigma(\|w\|),$
where the class $\mathcal{K}$ functions $\sigma_{h}$ and $\sigma_{w}$ exist due to uniform continuity of $\hat{h}_{\mathrm{PB}}$ and $f$ on the compact domain $\mathcal{D}_{\mathrm{PB}}$ as shown in [40]. The set $\mathcal{S}_{\mathrm{PB}}$ is 0-invariant with $\epsilon_{h}=0$ and $w=0$ and the set $\mathcal{D}_{\mathrm{PB}}$ is robust invariant, if, e.g., $\epsilon_{h}$ and $w$ is sufficiently small, i.e., it holds that $4\epsilon_{h}+\sigma(\|w\|)\leq\alpha_{f}\gamma_{f}-\bar{c}$, where $\bar{c}\coloneqq\max_{x\in\mathcal{D}_{\mathrm{PB}}}h_{\mathrm{PB}}(x)-c_{\Delta}\alpha_{3}\circ\alpha_{1}^{-1}(h_{\mathrm{PB}}(x))$, which ensures that $h_{\mathrm{PB}}(x^{+})\leq\alpha_{f}\gamma_{f}$ for all $x\in\mathcal{D}_{\mathrm{PB}}$.